OpenStudy (anonymous):

(x)(x+5)(3x-1)=7.6ft. Find the dimensions of the box in inches. Round to the nearest inch. x=width, x+5=length, and 3x-1=height.

OpenStudy (jlastino):

First you solve for x. Combine all the terms on the left side. What do you get?

OpenStudy (anonymous):

i know how to combine it, i need help on what to next after this: 3x^3+14x^2-5x

OpenStudy (anonymous):

Please help me! What do I set it equal to? 3x^3+14x^2-5x

OpenStudy (anonymous):

I get 3x^3+14x^2-5x... now what do i do?

OpenStudy (anonymous):

Excuse me? Do you know how to do this problem? Are u stuck? I need help! Plz say something!

OpenStudy (jlastino):

Sorry my internet connection is bad right now...i tried putting the equation here http://www.wolframalpha.com/input/?i=%28x%29%28x%2B5%29%283x-1%29%3D7.6

OpenStudy (anonymous):

but.... what about converting the feet into inches?

OpenStudy (jlastino):

just use the conversion factor 1 feet = 12 inches

OpenStudy (anonymous):

i know, but don't i get a whole different kind of answer?

OpenStudy (anonymous):

The dimensions in inches of a shipping box at We Ship 4 You can be expressed as width x, length x+5, and height 3x-1. The volume is about 7.6ft^3. Find the dimensions of the box in inches. Round to the nearest inch.

jhonyy9 (jhonyy9):

x(x+5)(3x-1)=7,6 so this is the volum x(x+5)=7,6/(3x-1) and this will be area of base

jhonyy9 (jhonyy9):

x^2 +5x =(7,6)/(3x-1) (3x-1)(x^2 +5x)=7,6 3x^3 +15x^2 -x^2 -5x =7,6 3x^3 +14x^2 -5x =7,6 3x^3 +14x^2 -5x -7,6 =0 multiplie both sides by 10 30x^3 +140x^2 -50x -76=0 divide both sides by 2 15x^3 +70x^2 -25x -38=0 so you need to solve this cubic equation and after you will get the x values so using these values of x you will can calcule easy the lenght and height of this box using these reports added in the first lines of this exercise width=x lenght=x+5 height=3x-1